Assignment 9

  1. If you run around a half-mile track twice in 20 minutes

a. What is the total distance traveled?

b. What is the speed in mi/hr (show work)?

c. What is the total displacement?

d. What is the velocity in mi/hr?

  1. You travel east 2 miles and then turn to the north and walk another 3 miles. What is your displacement (show work, make sure to show the direction as well)?
  2. You travel 6 miles at 40 degrees. How far north and how far east have you gone (show work)
  3. You travel 1400 miles at 65 mi/hr. How long does the trip take in hours and minutes (show work)?
  4. A sprinter goes from a standing position to 10 m/s in 2 seconds. What is her acceleration? How far has she traveled in those 2 seconds (show work)?
  5. The same sprinter slows down from 10 m/s to 5 m/s in 2 seconds. What is her acceleration? Why does the answer have the sign that it does (show work)?
  6. A truck decelerates from 15 m/s to 10 m/s in 200 meters. What is its average speed in m/s? How long does it take for the truck cover the 200 meters (show work)?
  7. An object is dropped from a height of 20 meters. Calculate how long it takes it to fall (show work).?
  8. A helicopter’s propeller is pushing the helicopter upwards with a force of 500 N. Gravity is applying a downward force of 300 N, and the wind is applying a downward force of 100 N. What is the net force on the helicopter (show your work)?
  1. An automobile is accelerating at 2 m/s2 to the east. The automobile’s mass is 1400 kg. What is the force on the automobile (show your work)?
  2. What does a 68 kg person weigh at sea level (show your work)?
  3. A 1400 kg automobile is going around a 50 meter radius curve at 4 m/s. What is the centripetal force on the automobile (show your work)?
  4. Two objects are separated by a distance of 2 meters. The first object has a mass of 50 kg, the second has a mass of 80 kg. What is the gravitation attraction between the two object (show your work)?
  1. A Buick rear-ends a stationary Honda. The Buick was traveling at 13 m/s at the moment of the inelastic collision. The Buick has a mass of 1800 kg, while the Honda has a mass of 1200 kg. What is the final velocity after the impact (show your work)?
  2. Using the data in Question 6, calculate the impulse and force of the two vehicles given a collision time of 0.1 seconds (show you work).

  1. Two objects collide in such a way that they act elastically. The first object is stationary and has a mass of 10 kg. The second object is moving at 2 m/s, and has a mass of 5 kg. What is the final velocity of the first object (show your work)?

  1. A weightlifter picks up a barbell with a force of 100 N a total distance of 2 meters. How much work is done (show your work)?
  1. Calculate the kinetic energy of a vehicle travelling at 25 m/s if the vehicle has a mass of 3000 pounds (show your work, must convert pounds into kilograms)?
  2. A worker lifts a barrel with a mass of 50 kg a distance of 2 meters with a force of 300 N in 10 seconds. Calculate the amount of work done and the worker’s power output in watts (show your work).
  3. Convert the number of watts you got for Question 12 into horsepower (show your work).
Answer #1

A. Total distance = 2 × 0.5 = 1 mile

B. Speed = distance / time

Here time = 20 min = 20/60 hr

Speed = 1/ 20/60 = 3 miles/ hr

C. Displacement = minimum distance between starring and ending point = 0

D. Velocity = displacement / time = 0